OpenStudy (erniellyd28):

1 3 + 2--- 3 ----------- 1 5 ----- - 1 6

Parth (parthkohli):

\[\rm \left({3 + 2{1 \over 3}}\right)\div \left(5{1 \over 6} - 1\right)\]?

OpenStudy (erniellyd28):

(3+6/3) divided by30/1 - 1. am i right so far?

Parth (parthkohli):

Hmm no.

Parth (parthkohli):

\[{2{1 \over 3}} = {2\cdot3+1 \over 3}\]\[{5{1 \over 6} = {5\cdot 6 + 1 \over 6}}\]

OpenStudy (erniellyd28):

3+7/3 divided by 30/1 - 1?

OpenStudy (erniellyd28):

so it would be 7/3 over 31over 6?

OpenStudy (erniellyd28):

or am i wrong?

OpenStudy (erniellyd28):

7/3 divided by 31/6?

OpenStudy (anonymous):

\[\rm \left({3 + 2{1 \over 3}}\right)\div \left(5{1 \over 6} - 1\right)\] \[= \rm \left({{9\over 3} + {7 \over 3}}\right)\div \left({31 \over 6} - {6 \over 6}\right)\]

OpenStudy (anonymous):

dont forget, when you divide two fractions you actualy multiply the first by the second flipped.